answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Where can one buy spinner rings?

Wal Mart, Target and Amazon offer spinner rings. Spinner rings can also be found at some jewelry stores or costume jewelry stores. Anxiety Spinner Rings seem to offer the best selection at reasonable prices.

Who do you calculate in a z-score for 120 plus 0.840 times 17.6?

120 + 0.84*17.6 is 134.784

The probability associated with a z-score of 3 is 0.0013

By z = 6, it drops to 1 billionth.

A z-score of 10 or more is near enough 0 to make no difference.

Are you sure about the question? Also, whatever the question, you do not calculate someone!

If rolling one fair die what is the probability of odd or less than 5?

The odd values are 1, 3, 5 and less than 5 values are 1, 2, 3, 4. The probability is the number of ways success can occur divided by the sample space. The sample space is 6. The values that we can have to fir the "success" definition is 5. So, probability of odd or less than 5 is 5/6.

Why is time a more important factor in determinig probable cause to search that it is in determining probable cause to arrest?

When a search for evidence is contemplated, the search needs to be done in a timely manner before the evidence is destroyed, disposed of, or moved. The state typically has years to accumulate enough probable cause to arrest someone for a crime.

What is the rule that applies to the following sentence 0 divieded by 70?

The rule that applies is USE YOUR SPELL CHECKER. There is no such word as divieded.

Why you need sampling distribution?

in order to calculate the mean of the sample's mean and also to calculate the standard deviation of the sample's

What is certain event in probability?

A certain event is any event that will definitely happen.

For example, the sunrise tomorrow is something that is certain to happen (although it might be obscured from view by rain, snow, fog or something else); this will certainly happen.

How many queens in adeck of 52 cards?

4 queens.

In a deck of 52 cards there are only ever 4 of any card. 1 heart, 1 spade, 1 club, 1 diamond. (eg. 1 queen of hearts, 1 queen of spades, 1 queen of clubs, 1 queen of diamonds)

What does sku mean?

It stands for Stock Keeping Unit.The SKU number is a VITAL piece of information to every store. Combined with the UPC number, the SKU identifies the product within the companies inventory database. It not only supplies information about the items cost, but also about the number of items in the store, the number on back order and the number that can be purchased from other stores in the chain. Without the SKU and UPC numbers, vendors would have no idea the amount of items they carry, or what they are spending or earning on that product and their profit margin would be impossible to track.

Can a yes or no survey be done on a ordinal scale?

no as ordinal is in a precise order and you casnnot predict peoples answers to the questions as you would have to have all of the yes' then all of the nos.

A bag contain 4 balls 2 balls are drawn at random and are found to be white What is the probability that all balls are white?

it has three cases 1st E1- two balls are white
E2- three balls are white
E4-four balls are white
E4/w= (p(E3)*4C2/4C2)/(P(E1)*2C2/4C2+P(E2)3C2/4C2+P(E3)*4C2/4C2)
WHERE-P(E1)=P(E2)=P(E3)=1/3
----------------------------------------------------------------------------------------------------

2nd opinion

Let's say we have 3 boxes with 4 balls each.

Box A has 4 white balls.

Box B has 3 white balls.

Box C has 2 white balls.

The probability of drawing 2 W balls from;

Box A, P(2W│A)=

(4/4)∙(3/3)=

1

Box B, P(2W│B)=

(3/4)∙(2/3)=

1/2

Box C, P(2W│C)=

(2/4)∙(1/3)=

1/6

Say the probability of picking any of the 3 boxes is the same, we have;

P(A)=

1/3

P(B)=

1/3

P(C)=

1/3

Question is, given the event of drawing 2 W balls from a box taken blindly

from the 3 choices, what is the probability that the balls came from box A,

P(A│2W).

Recurring to Bayes Theorem:

P(A│2W)=

[P(A)P(2W│A)]/[P(A)P(2W│A)+P(B)P(2W│B)+P(C)P(2W│C)]=

[(1/3)(1)]/[(1/3)(1)+(1/3)(1/2)+(1/3)(1/6)]=

6/10=

0.60=

60%

P(A│2W)=

0.60=

60%



Read more:Solution_to_a_bag_contains_4_balls_Two_balls_are_drawn_at_random_and_are_found_to_be_white_What_is_the_probability_that_all_balls_are_white

If a coin and a dice is tossed what is the probaility of getting a heads and six?

Heads comes up half the time, sis one time in six tosses so the combined odds of heads and six is 1/2 x 1/6 = 1/12, one time in 12.

Can a woman still be menstruating when pregnant?

No. If you are pregnant and bleeding you should contact your Dr right away.

Two cards are drawn from a deck of cards. what is the probability that both cards are spades?

This question is a little bit tricky. In a deck of 52 cards, one-fourth or 13 cards are spades. So, the chance of drawing one spade = 13/52 or 0.25. If a second card drawn, there's one less spade in the deck, so the probability on the second draw is 12/51.

The probability of drawing two spades from a deck is 0.25 x 12/51 = 0.058824 This is called sampling without replacement. In quality control, it is very common to sample without replacement as bad parts are discarded.

If we consider drawing one card, putting it back in the deck, regardless if it is a spade or not, then reshuffling the deck and drawing the second card, the probability is 0.25 x 0.25 = 0.0625, a bit higher with replacement. This is the same as 1/4 x 1/4 = 1/8 or saying the odds are 1:8.

I've included a couple of links on sampling with replacement and without replacement. Generally, for calculating statistics, we attempt to get independent results. The draw of one card, will reduce the population, and change the probabilities on the second draw, so sampling without replacement is not independent sampling.

See related links.

How many three of diamonds are in a 52 deck of cards?

There is 1 Three of Diamonds in a standards 52 card deck of cards.

What is the probability that 12 rolls of a die will show exactly three fours?

When a die is rolled once, the probability of a 4 showing up is 1/6.

Apply the binomial probability for finding the probability of exactly three fours out of 12 throws of a die.

n=12 (number of throws)

p=1/6 (probability of a four in a single throw)

x = 3 (number of times out of 12 , a four showing up)

P(x=3) = 12C3 (1/6)^3 (5/6)^(12-3)

= 12C3 (1/6)^3 (5/6)^9 = 0.197443